# Magnetic signals from oceanic tides: new satellite observations and applications

### v1

@@ Abstract
Tidal flow of seawater across the Earth’s magnetic field induces electric currents and magnetic fields within the ocean and solid Earth. The amplitude and phase of the induced fields depends on electrical properties of both the seawater and the solid Earth, thus can be used as a proxy to study seabed properties or potentially for monitoring long-term trends in the global ocean climatology. This paper presents new global oceanic tidal magnetic field models and their uncertainties for four tidal constituents, including $M_{2},N_{2},O_{1}$ and for the first time $Q_{1}$. Models are obtained through a robust least-squares analysis of magnetic field observations from the Swarm and CHAMP satellites using a specially designed data selection scheme. We compare the retrieved magnetic signals with several alternative models reported in the literature. Additionally, we validate them using a series of high-resolution global 3-D electromagnetic simulations and place constraints on the conductivity of sub-oceanic mantle for all tidal constituents, revealing an excellent agreement between all tidal constituents and the oceanic upper mantle structure.
@@ 1 Introduction
The tidal forcing contributes energy at specific frequencies, referred to as tidal harmonic constituents (for brevity, we will use this term for all spectral lines where tidal gravitational forcing inputs energy), which can be accurately predicted from the development of the tide-generating potential [3, 4]. Tidal constituents are most often distinguished based on their frequency (e.g. quarter-diurnal, semi-diurnal, diurnal and long-period) and/or origin (e.g. lunar, solar or nonlinear/mixed tides). The strongest tidal constituents and some of their characteristics are listed in Table 1. Note that effects associated with the bathymetry, ocean stratification and other hydrodynamic effects result in the redistribution of the energy and local variations in height and velocity, but they do not alter harmonic frequencies, making tides highly predictable [5]. A more refined and complete description of oceanic tides and their properties can be found in the dedicated literature.
@@
$\mathrm{K}_{2}^{\mathrm{L}}$ 11.967 $2\omega_{L}+2\omega_{1}(=2\Omega)$ declinational to $\mathbf{M}_{2}$ 0.0768
@@
The atmosphere (including the ionosphere) is also subject to gravitational tides. However, in addition to the gravity forces, the ionosphere experiences daily thermal tides generated by solar heating. These convective motions of the ionosphere dominate the motions generated by gravitational forces. As a result, motionally induced electric currents due to gravitational tides are dwarfed by the currents generated by thermal winds. Additionally, while the electrical conductivity of the ocean remains virtually constant on diurnal time scales, the conductivity structure of the ionosphere varies strongly with the local time and depends on upstream solar wind conditions [7]. In addition, the ionospheric conductivity is highly anistropic. These considerations are important when considering the design of optimal data selection schemes and the validation of results, particularly for tidal constituents with periods very close or equal to 24 and 12 hours.
@@
Early works in the terrestrial magnetism also recognized solar and lunar components in the daily variations of the geomagnetic field (see a discussion in [9]). These variations were mostly ascribed to the ionospheric dynamo, and a smaller internal component of the observed diurnal variations was explained by electromagnetic induction effects in the solid Earth [10]. A comprehensive review of earlier works dedicated to studying the geomagnetic tides of external origin is presented in [11]. The presence of the intrinsic ocean-generated component in observed geomagnetic tides was discussed by W. van Bemmelen in the work titled ”Die lunare Variation des Erdmagnetismus” published in 1912 (the original study [12] appears not to be available online, a brief summary can be retrieved from the in memoriam [13]. Decades later, a systematic analysis of coastal, inland and first ocean-bottom observations [14, 15, 16, 17], as well as development of an analytic theory describing the responsible variations [18, 19, 20] allowed many groups to model oceanic tidal magnetic signals and validate the growing body of observations. In the majority of early geomagnetic field studies, ocean-induced magnetic fields were seen as noise or a nuisance component rather than the signal of interest [16]. Therefore, little attention was paid to this part of the observed signal beyond proving its existence. Nevertheless, some early works did recognize the potential of using oceanic tidal magnetic fields for studying the electric properties of the seabed, upper mantle or even seawater currents [8, 19, 21].
@@
A major upsurge in the interest of the geophysical community came with the satellite era. The magnetic signature of a principal lunar $M_{2}$ tide was isolated in satellite magnetic field observations a few years into the satellite mission CHAMP [22]. Later, a CHAMP-based comprehensive geomagnetic field model CM5 [23] used a potential field representation of the $M_{2}$ tidal magnetic signal and delivered a global model of the $M_{2}$ tide magnetic signal (data from earlier missions SAC-C and Ørsted were also included, although they play a less significant role in characterizing oceanic magnetic signals due to higher altitudes). This model was accurate enough to constrain the global electrical conductivity of the sub-oceanic upper mantle for the first time [24, 25]. The science program of the ESA Swarm mission, launched in November 2013, defined magnetic signals due to oceanic tides as a mission science objective [26] and magnetic signals due to the dominant $M_{2}$ tide have indeed been retrieved from the Swarm satellite data [27, 28]. The work of [29] combined both CHAMP and Swarm data and, in addition to the $M_{2}$ constituent, also added $N_{2}$ and $O_{1}$ tides. Subsequent works by other groups also succeeded in extracting $M_{2},N_{2}$, and $O_{1}$ components [30, 31]. All these studies adopted a potential field representation for the tidal magnetic signals as in [23] and used vector, scalar and approximate gradients (along-track differences for the CHAMP and both along-track as well as East-West differences between Alpha and Charlie satellites for the Swarm constellation) data. On the other hand, these studies differ in the details of the data selection and geomagnetic field modelling approaches. In short, our previous work [29] used the so-called sequential approach, where the magnetic field residuals used for the global tidal harmonic analysis were obtained by subtracting the CHAOS-6 model of core and magnetospheric fields [32]. This approach will be further developed in the present study. In contrast, the works of [30] and [31] used a Comprehensive Inversion scheme handling all sources simultaneously and a method based on a Kalman filter, respectively.
@@
In this study we will address the problem of mapping the magnetic signals due to oceanic tides with satellite observations, characterize the uncertainty of the obtained models and present a systematic comparisons with detailed 3-D electromagnetic simulations and other models reported in the literature. Finally, we will use new tidal magnetic field models to infer constraints on the electrical conductivity of the lithosphere-asthenosphere system in the sub-oceanic mantle.
@@ 2.1 Governing Equations
Electric currents and magnetic fields induced by oceanic tides are low-frequency phenomena. Therefore, their generation and evolution can be accurately described by adopting a magneto-quasi static approximation to the Maxwell’s equations [33] and assuming a linear non-polarized medium without additional magnetization currents:
@@
$\begin{split}\vec{J}&=\sigma\left[\vec{E}+\vec{v}\times\vec{B}\right]\\ &\triangleq\sigma\left[\vec{E}+\vec{v}\times\vec{B}_{c}\right]\end{split}$ (2)
@@
2.2 Model parameterization
@@
By calculating frequencies $\omega$ from the periods given in Table 1, SH coefficients for different harmonic tidal constituents were estimated up to a maximum spherical harmonic degree $N_{\textnormal{max}}$. The values for $N_{\textnormal{max}}$ were chosen experimentally for each constituent to ensure reasonable signal-to-noise levels while retaining the highest possible spatial resolution. A Huber-weighted robust least-squares method was used to retrieve the SH coefficients $\tau$ (the method is documented in our previous study [29]). Assuming that $\mathbf{A}$ is a matrix relating SH coefficients to the observations and $\mathbf{W}$ is a diagonal matrix containing Huber weights, the model covariance matrix $\mathbf{C}_{\tau}=\left(\gamma^{-2}\mathbf{A}^{T}\mathbf{W}\mathbf{A}\right)^{-1}$ was calculated for each tidal constituent, with $\gamma=2$ nT being the assumed prior data standard deviation, consistent with CHAOS-7 residual statistics (cf. Figure 6 in [37]) for CHAMP and Swarm vector field components. $\delta\mathbf{\tau}=\mathrm{diag}(C_{m})^{1/2}$ are then the model standard deviations, which we use as a measure of the model uncertainty propagating it to the predicted magnetic fields.
2.3 Numerical simulations
In subsequent sections, we will use numerical simulations of ocean tidal induced magnetic fields from different tidal constituents for two purposes: (i) to validate the extracted signals and (ii) to test sensitivity of the extracted signals to the conductivity contrast across the lithosphere-asthenosphere boundary. Since tidal magnetic signals are strongly affected by lateral electrical conductivity contrasts [38], a realistic 3-D model of the ocean and marine sediments conductivity is required. Accordingly, a full 3-D EM induction solver is needed to account for all inductive and galvanic effects for a complex tidal electrical current density and a 3-D electrical conductivity distribution. In this study, we rely on a novel 3-D EM code [39] that is based on the adaptive Finite Element Method implemented within the deal.II library [40] and solves eqs. (1) numerically in a sphere using an optimal multi-grid solver [41]. A hexahedral locally refined mesh was used and the solution is approximated by second-order $curl$-confirming Nédélec basis functions [42] within each grid cell. This results in uniform horizontal resolution of $\approx 10$ km at the surface and at satellite altitude. This is the reference solution for task (i) listed above. Given the limited computational resources, a lower resolution grid was used for the sensitivity study (ii) where many forward solutions are required.
For the electrical conductivity of the ocean and marine sediments, we used the realistic 3-D model of [43], which relies on validated observations of ocean climatology from the World Ocean Atlas and the thermodynamic equation of state for seawater (TEOS-10). For land areas, the model of [44] was taken. The underlying mantle section is assigned a new radial reference conductivity profile derived from the joint inversion of recent Swarm-based magnetospheric and $M_{2}$ tidal magnetic signals (specifically, Comprehensive Inversion Swarm products MMA_0906 and MTI_0901) following the methodology described in our earlier work [45]. The TPXO-9 model was used for the velocities of the tidal constituents [46]. The magnetic field $\vec{B}_{c}$ was calculated using the IGRF-11 model evaluated at 1st November 2014.
@@ 3 Data
Figure 1: Number of observations used in the analysis as a function of time. Scalar and vector observations are denoted by $F$ and $\vec{B}$, respectively. The corresponding scalar and vector magnetic field gradients are denoted $\delta F$ and $\delta\vec{B}$.
In order to focus on the oceanic tidal signal the time-dependent internal (core) field and magnetospheric fields from the CHAOS-7.12 model [37], and the lithospheric field from the LSC-1 field model [50] were removed from the observations; only the residuals after these corrections were used for constructing the field models described below.
@@ 4.1 Tidal magnetic signals
Figure 2: Maps of the real and imaginary parts of the observed, $B_{r}^{\mathrm{obs}}$, and numerically simulated, $B_{r}^{\mathrm{mod}}$, ocean tidal radial magnetic fields due to the semi-diurnal $M_{2}$ and $N_{2}$ tidal constituents. Fields are given at the altitude of 430 km.
@@
Figures 2-3 show maps of the radial magnetic field components, at satellite altitude, derived in this study. Both observed and modelled fields are shown and can be compared. Compared to our previous study, GO19, significant improvements in signals at high latitudes are achieved. For instance, strong tidal signals in the Greenland-Iceland region (most visible in $M_{2}$) are now reliably retrieved and exhibit a good agreement with simulations. Contamination at high latitudes is more apparent in the magnetic fields retrieved for the $N_{2}$ and $O_{1}$ tides, while the agreement at mid and equatorial latitudes is generally excellent. Even the weak $Q_{1}$ signal exhibits a good degree of similarity between the observed and simulated signals in terms of the general patterns, although the spatial resolution and signal-to-noise ratio remain low. With $N_{\mathrm{max}}=32$ for the $M_{2}$ signal we are approaching a spatial resolution of $1000-2000$ km. Regional features caused by variations in the tidal transport, e.g. around islands, can now be retrieved from data and validated by simulations. Figure 2 highlights a few such regions with strong local gradients in the tidal magnetic fields that are visible in both observations and simulations.
@@
Figure 4: Standard deviation (in nT) maps calculated from the model covariance matrix $\mathbf{C}_{m}$ for each tidal constituent and propagated to the radial magnetic field at 430 km altitude.
@@
$\mathrm{RMSD}(r)=\sqrt{\sum{\frac{\left[\delta B_{r}(\vec{r})\delta B_{r}(\vec{r})^{*}\right]\mathrm{d}S}{\mathrm{d}S}}},$ (4)
where the sum runs over all surface elements (we used a 1 resolution, yielding 64,800 surface elements), star denotes the complex conjugation, $\mathrm{d}S=\sin{\theta}\Delta\phi\Delta\theta$ is the surface element area and $\delta B_{r}(\vec{r})=B_{r}^{\mathrm{obs}}(\vec{r})-B_{r}^{\mathrm{mod}}(\vec{r})$. It is necessary to take into account the variable physical area of spherical surface elements to avoid a bias in polar regions.
Table 2: Left: Root-mean-square deviations (RMSD) between observed tidal signals and 3-D numerical simulation for four considered tidal constituents. Radial magnetic field component at the altitude of 430 km was used. Right: corresponding maximum SH degree used to calculate the magnetic fields.
@@
2 Although higher degree expansions were reported in the original study [30], the authors stated that signals up to the herein specified degrees are reliable.
Table 2 lists RMSD values for all the compared models. The use of a numerical simulation for the validation has several implications. It is an independent result that relies on the actual equations that simulate all relevant physics. Furthermore, by using water transport from the satellite altimetry-based TPXO-9 model for the electric current term we ensure that simulated motionally induced signals are generated by realistic tidal currents. On the other hand, a key unknown in eqns (1)-(2) is the electrical conductivity of the underlying crust and upper mantle. For this we used a global average radial conductivity profile for the upper mantle derived from the CI-based $M_{2}$ signal, hence the simulated signals will naturally tend towards the CI9 model. Although this creates somewhat favourable conditions for the CI9 tidal field model, it is the GFO24 model that exhibits the smallest RMSD. We stress that tidal magnetic signals are also sensitive to lateral conductivity variations in the crust and mantle [52], hence non-zero RMSD values can be ascribed to several factors: (i) lateral conductivity variations in the subsurface not accounted for by the present numerical simulation, (ii) errors in the representation of the current generating terms (water transport, core field, seawater conductivity) and (iii) numerical/discretization errors. The last two factors were previously quantified and should not generally exceed $\sim$0.1 nT for the strongest $M_{2}$ tide [25, 38].
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Pearson correlation coefficients between different models of the observed signals and a 3-D discrete numerical simulation (denoted DNS) for the four tidal constituents. Correlations were calculated between real parts of the radial magnetic field maps at the altitude of 430 km (shown in Figures 2-3). Correlation coefficients for imaginary parts are similar.
@@
Although the structure of spectra are generally similar, there are some differences: SH degrees $\geq 15-20$ appear to be damped in the KALMAG model, which prevents comparison with other models at these degrees. The new GFO24 models exhibit better agreement with simulations at all scales compared to the predecessor study GO19. In particular, the low degree SH coefficients are closer to the 3-D simulations, highlighting a reduced contamination by external field sources. We stress that the presented SH power spectra were calculated at Earth’s surface. Such downward continuation is known to amplify noise [53]; the generally good agreement, even for smaller $N_{2}$ and $O_{1}$ constituents, is a remarkable achievement of modern satellite magnetometry. That said, we do observe a larger misfit at higher SH degrees for the GFO24-HR models compared to the GFO24 models where they overlap, indicating higher variances at high SH degrees in the GFO24-HR models. This stems from a fundamental trade-off between the resolution and variance, which is inherent to the geomagnetic field inverse problem solved here.
Figure 6: The $R_{n}$ power spectra (eq. 5) of the time-averaged ocean tidal magnetic fields due to four tidal constituents for different models. Black lines denote 3-D numerical simulation results. Spectra were calculated for the reference surface radius $R_{E}=6371.2$ km.

### v4

@@ Abstract
Tidal flow of seawater across the Earth’s magnetic field induces electric currents and magnetic fields within the ocean and solid Earth. The amplitude and phase of the induced fields depends on electrical properties of both the seawater and the solid Earth, thus can be used as a proxy to study seabed properties or potentially for monitoring long-term trends in the global ocean climatology. This paper presents new global oceanic tidal magnetic field models and their uncertainties for four tidal constituents, including $M_{2},N_{2},O_{1}$ and $Q_{1}$, which was not reliably retrieved previously. Models are obtained through a robust least-squares analysis of magnetic field observations from the Swarm and CHAMP satellites using a specially designed data selection scheme. We compare the retrieved magnetic signals with several alternative models reported in the literature. Additionally, we validate them using a series of high-resolution global 3-D electromagnetic simulations and place constraints on the conductivity of sub-oceanic mantle for all tidal constituents, revealing an excellent agreement between all tidal constituents and the oceanic upper mantle structure.
@@ 1 Introduction
The tidal forcing contributes energy at specific frequencies, referred to as tidal harmonic constituents (for brevity, we will use this term for all spectral lines where tidal gravitational forcing inputs energy), which can be accurately predicted from the development of the tide-generating potential [3, 4]. Tidal constituents are most often distinguished based on their frequency (e.g. quarter-diurnal, semi-diurnal, diurnal and long-period) and/or origin (e.g. lunar, solar or nonlinear/mixed tides). The strongest tidal constituents and some of their characteristics are listed in Table 1. Note that effects associated with the bathymetry, ocean stratification and other hydrodynamic effects result in the redistribution of the energy and local variations in height and velocity, but they do not alter harmonic frequencies, making tides highly predictable. A more refined and complete description of oceanic tides and their properties can be found in the dedicated literature [5, 6].
@@
$\mathrm{K}_{2}^{\mathrm{L}}$ 11.967 $2\omega_{L}+2\omega_{1}(=2\Omega)$ declinational to $\mathrm{M}_{2}$ 0.0768
@@
The atmosphere (including the ionosphere) is also subject to gravitational tides. However, in addition to the gravity forces, the ionosphere experiences daily thermal tides generated by solar heating. These convective motions of the ionosphere dominate the motions generated by gravitational forces. As a result, motionally induced electric currents due to gravitational tides are dwarfed by the currents generated by thermal winds. Additionally, the electrical conductivity of the ocean remains virtually constant on diurnal time scales, with the exception of the Sea Surface Temperature and Salinity [8, 9], although these near-surface variations have a negligible effect on tidal magnetic signals that are primarily controlled by the depth-integrated seawater conductance [10]. In contrast, the conductivity structure of the ionosphere varies strongly with the local time and depends on upstream solar wind conditions [11]. In addition, the ionospheric conductivity is highly anistropic. These aspects are important when considering the design of optimal data selection schemes and the validation of results, particularly for tidal constituents with periods very close or equal to 24 and 12 hours.
@@
Early works in terrestrial magnetism also recognized solar and lunar components in the daily variations of the geomagnetic field (see a discussion in [13]). These variations were mostly ascribed to the ionospheric dynamo, and a smaller internal component of the observed diurnal variations was explained by electromagnetic induction effects in the solid Earth [14]. A comprehensive review of earlier works dedicated to studying the geomagnetic tides of external origin is presented in [15]. The presence of the intrinsic ocean-generated component in observed geomagnetic tides was discussed by W. van Bemmelen in the work titled ”Die lunare Variation des Erdmagnetismus” published in 1912 (the original study [16] appears not to be available online, a brief summary can be retrieved from the in memoriam [17]). Decades later, a systematic analysis of coastal, inland and the first ocean-bottom observations [18, 19, 20, 21], as well as development of an analytic theory describing the responsible variations [22, 23, 24] allowed many groups to model oceanic tidal magnetic signals and validate the growing body of observations. In the majority of early geomagnetic field studies, ocean-induced magnetic fields were seen as noise or a nuisance component rather than the signal of interest [20]. Therefore, little attention was paid to this part of the observed signal beyond proving its existence. Nevertheless, some early works did recognize the potential of using oceanic tidal magnetic fields for studying the electric properties of the seabed, upper mantle or even seawater currents [12, 23, 25].
@@
A major upsurge in the interest of the geophysical community came with the satellite era. The magnetic signature of a principal lunar $M_{2}$ tide was isolated in satellite magnetic field observations a few years into the satellite mission CHAMP [26]. Later, a CHAMP-based comprehensive geomagnetic field model CM5 [27] used a potential field representation of the $M_{2}$ tidal magnetic signal and delivered a global model of the $M_{2}$ tide magnetic signal (data from earlier missions SAC-C and Ørsted were also included, although they play a less significant role in characterizing oceanic magnetic signals due to a weaker signal at their higher altitudes). This model was accurate enough to constrain the global electrical conductivity of the sub-oceanic upper mantle for the first time [28, 29]. The science program of the ESA Swarm mission, launched in November 2013, defined magnetic signals due to oceanic tides as a mission science objective [30] and magnetic signals due to the dominant $M_{2}$ tide have indeed been retrieved from the Swarm satellite data [31, 32]. The work of [33] combined both CHAMP and Swarm data and, in addition to the $M_{2}$ constituent, also added $N_{2}$ and $O_{1}$ tides. Subsequent works by other groups also succeeded in extracting $M_{2},N_{2}$, and $O_{1}$ components [34, 35]. All these studies adopted a potential field representation for the tidal magnetic signals as in [27] and used vector, scalar and approximate gradients (along-track differences for the CHAMP and both along-track as well as East-West differences between Alpha and Charlie satellites for the Swarm constellation) data. On the other hand, these studies differ in the details of the data selection and geomagnetic field modelling approaches. In short, our previous work [33] used the so-called sequential approach, where the magnetic field residuals used for the global tidal harmonic analysis were obtained by subtracting models of core and magnetospheric fields as given by the CHAOS-6 model [36]. This approach will be further developed in the present study. In contrast, the works of [34] and [35] used a Comprehensive Inversion scheme handling all sources simultaneously and a method based on a Kalman filter, respectively.
@@
In this study we will address the problem of mapping the magnetic signals due to oceanic tides with satellite observations, characterize the uncertainty of the obtained models and present systematic comparisons with detailed 3-D electromagnetic simulations and other models reported in the literature. Finally, we will use new tidal magnetic field models to infer constraints on the electrical conductivity of the lithosphere-asthenosphere system in the sub-oceanic mantle.
@@ 2.1 Governing Equations
Electric currents and magnetic fields induced by oceanic tides occur at frequencies where displacement currents are negligible [37]. Therefore, generation and evolution of tidal electromagnetic fields can be accurately described by adopting a magneto-quasi static approximation to Maxwell’s equations [37] and assuming a linear non-polarized medium without additional magnetization currents:
@@
$\begin{split}\vec{J}&=\sigma\left[\vec{E}+\vec{v}\times\vec{B}\right]\\ &\approx\sigma\left[\vec{E}+\vec{v}\times\vec{B}_{c}\right]\end{split}$ (2)
@@
2.2 3-D Numerical Simulations
A direct numerical simulation (DNS), hereinafter denoted by $\vec{B}^{\mathrm{DNS}}(\vec{r})$, can be obtained by solving eqs. 1-2 numerically. This section provides details on how DNS is performed in this study and when it is employed.
In subsequent sections, we will use DNS of ocean tidal induced magnetic fields from different tidal constituents for two purposes: (i) to validate the observed signals and (ii) to test sensitivity of the observed signals to the conductivity contrast across the lithosphere-asthenosphere boundary. Since tidal magnetic signals are strongly affected by lateral electrical conductivity contrasts [40], a realistic 3-D model of the ocean and marine sediments conductivity is required. Accordingly, a full 3-D EM induction solver is needed to account for all inductive and galvanic effects for a complex tidal electrical current density and a 3-D electrical conductivity distribution. In this study, we rely on a novel 3-D EM code [41] that is based on the adaptive Finite Element Method implemented using the deal.II library [42] and solves eqs. (1) numerically in a sphere using an optimal multi-grid solver [43]. A hexahedral locally refined mesh was used and the solution is approximated by second-order $curl$-confirming Nédélec basis functions [44] within each grid cell. This results in uniform horizontal resolution of $\approx 10$ km at the surface and at satellite altitude. This is the reference solution for task (i) listed above. Given the limited computational resources, a lower resolution grid was used for the sensitivity study (ii) where many forward solutions are required.
For the electrical conductivity of the ocean and marine sediments, we used the realistic 3-D model of [45], which relies on validated observations of ocean climatology from the World Ocean Atlas and the thermodynamic equation of state for seawater (TEOS-10). In future, this can be updated with a validated product [46, 47]. For land areas, the model of [48] was taken. The underlying mantle section is assigned a new radial reference conductivity profile derived from the joint inversion of recent Swarm-based magnetospheric and $M_{2}$ tidal magnetic signals (specifically, Comprehensive Inversion Swarm products MMA_0906 and MTI_0901) following the methodology described in our earlier work [49]. The TPXO-9 model was used for the velocities of the tidal constituents [50]. The magnetic field $\vec{B}_{c}$ was calculated using the IGRF-11 model evaluated at 1st November 2014.
2.3 Observational Models
@@
By calculating frequencies $\omega$ from the periods given in Table 1, SH coefficients for each harmonic tidal constituent were estimated up to a maximum spherical harmonic degree $N_{\textnormal{max}}$ from pre-processed satellite geomagnetic field observations (see Section 3). The values for $N_{\textnormal{max}}$ were chosen experimentally for each constituent to ensure reasonable signal-to-noise levels while retaining the highest possible spatial resolution. A Huber-weighted robust least-squares method was used to retrieve the SH coefficients $\tau$ (the method is documented in our previous study [33]). Hereinafter, the radial magnetic field due to a given tidal constituent at a location $\vec{r}$ above the ground, synthesized from corresponding estimated SH coefficients using eq. (3), will be denoted as $B_{r}^{\mathrm{obs}}(\vec{r})$.
To quantify errors of a SH model given by coefficients $\tau$, we perform a simple error propagation. Assuming that $\mathbf{A}$ is a matrix relating SH coefficients to the observations for a given tidal constituent and $\mathbf{W}$ is a diagonal matrix containing Huber weights, the model covariance matrix $\mathbf{C}_{\tau}=\left(\gamma^{-2}\mathbf{A}^{T}\mathbf{W}\mathbf{A}\right)^{-1}$ was calculated for each tidal constituent, with $\gamma=2$ nT being the assumed prior data standard deviation, consistent with CHAOS-7 residual statistics (cf. Figure 6 in [52]) for CHAMP and Swarm vector field components. $\delta\mathbf{\tau}=\mathrm{diag}(C_{\tau})^{1/2}$ are then the model standard deviations, which we use as a measure of the model uncertainty propagating it to magnetic fields. Since Huber weights and periods are different for different constituents, this procedure is repeated for each tidal constituent independently.
@@ 3 Data
Figure 1: Number of observations used in the analysis as a function of time. Scalar and vector observations are denoted by $F$ and $\vec{B}$, respectively. The corresponding scalar and vector magnetic field gradients are denoted $\delta F$ and $\delta\vec{B}$. Each component of a vector field counts as a single datum.
In order to focus on the oceanic tidal signal, the time-dependent internal (core) field and magnetospheric fields from the CHAOS-7.12 model [52], and the lithospheric field from the LSC-1 field model [58] were removed from the observations; only the residuals after these corrections were used for constructing the field models described below.
Compared with the CHAOS-6 model used in our previous study [33], the CHAOS-7 model employed during the data processing here involves a number of improvements. The basic field model parameterization is similar but it was constructed using updated datasets from the three Swarm satellites and ground observatories, as well as a new source of vector magnetic field data from the Cryosat-2 satellite. CHAOS-7 also uses stricter geomagnetic quiet time data selection criteria compared with CHAOS-6, relaxed temporal regularization of the high degree internal field and uses the diagonal part a 3D Q-response matrix [59] when induced fields are computed. Full details of the CHAOS-7 model can be found in the reference publication [52]; CHAOS-7.12 was the most up to data version of the model available at the time of data processing, it was 12th update of the CHAOS-7 model made using the Swarm and ground observatory data as available in September 2022.
@@ 4.1 Tidal magnetic signals
Figure 2: Maps of the real and imaginary parts of the observed, $B_{r}^{\mathrm{obs}}$, as given by the GFO24-HR model, and numerically simulated, $B_{r}^{\mathrm{DNS}}$, ocean tidal radial magnetic fields due to the semi-diurnal $M_{2}$ and $N_{2}$ tidal constituents. Fields are given at the altitude of 430 km. Red rectangles highlight the regions discussed in detail in the text.
@@
Figures 2-3 show maps of the radial magnetic field components, at satellite altitude, derived in this study. Both observed and modelled fields are shown and can be compared. Compared to our previous study, GO19, significant improvements in signals at high latitudes are achieved. For instance, strong tidal signals in the Greenland-Iceland region (most visible in $M_{2}$) are now reliably retrieved and exhibit a good agreement with simulations. However, one clear sign of contamination is significant magnetic signals far from the oceans, deep inland of the continents, mostly caused by ionospheric currents. Contamination at high latitudes is more apparent in the magnetic fields retrieved for the $N_{2}$ and $O_{1}$ tides, while the agreement at mid and equatorial latitudes is generally excellent. Even the weak $Q_{1}$ signal exhibits a good degree of similarity between the observed and simulated signals in terms of the general patterns, although the spatial resolution and signal-to-noise ratio remain low. With $N_{\mathrm{max}}=32$ for the $M_{2}$ signal our models approach a spatial resolution of $1000-2000$ km at the Earth’s surface. Regional features caused by variations in the tidal transport, e.g. around islands, can now be retrieved from data and validated by simulations. Figure 2 highlights a few such regions with strong local gradients in the tidal magnetic fields that are visible in both observations and simulations. For instance a region southeast of Africa with regional amplifications (south of the Madagascar) and weakening of the field (around French Southern and Antarctic Lands) are well characterized.
@@
Figure 4: Standard deviation (in nT) maps calculated from the model covariance matrix $\mathbf{C}_{\tau}$ for each tidal constituent and propagated to the radial magnetic field at 430 km altitude.
@@
$\mathrm{RMSD}(r)=\sqrt{\frac{\sum{\left[\delta B_{r}(\vec{r})\delta B_{r}(\vec{r})^{*}\right]\mathrm{d}S}}{\sum{\mathrm{d}S}}},$ (4)
where the sum runs over all surface elements (we used a 1 resolution, yielding 64,800 surface elements), star denotes the complex conjugation, $\mathrm{d}S=\sin{\theta}\Delta\phi\Delta\theta$ is the surface element area and $\delta B_{r}(\vec{r})=B_{r}^{\mathrm{obs}}(\vec{r})-B_{r}^{\mathrm{DNS}}(\vec{r})$. It is necessary to take into account the variable physical area of spherical surface elements to avoid a bias in polar regions.
Table 2: Left: Root-mean-square deviations (RMSD) between observed tidal signals and 3-D DNS for four considered tidal constituents (calculated as described in Section 2.2). Radial magnetic field component at the altitude of 430 km was used. Right: corresponding maximum SH degree used to calculate the magnetic fields.
@@
2 Although higher degree expansions were reported in the original study [34], the authors stated that only signals up to the herein specified degrees are reliable.
Table 2 lists RMSD values for all the compared models. The use of a numerical simulation for the validation has several implications. It is an independent result that relies on the actual equations that simulate all relevant physics. Furthermore, by using water transport from the satellite altimetry-based TPXO-9 model for the electric current term we ensure that simulated motionally induced signals are generated by realistic tidal currents. On the other hand, a key unknown in eqns (1)-(2) is the electrical conductivity of the underlying crust and upper mantle. For this we used a global average radial conductivity profile for the upper mantle derived from the CI-based $M_{2}$ signal, hence the simulated signals will naturally tend to be consistent with the CI9 model. Although this creates somewhat favourable conditions for the CI9 tidal field model, it is the GFO24 model that exhibits the smallest RMSD. We stress that tidal magnetic signals are also sensitive to lateral conductivity variations in the crust and mantle [61], hence non-zero RMSD values can be ascribed to several factors: (i) lateral conductivity variations in the subsurface not accounted for by the present numerical simulation, (ii) errors in the representation of the current generating terms (water transport, core field, seawater conductivity) and (iii) numerical/discretization errors. The last two factors were previously quantified and should not generally exceed $\sim$0.1 nT for the strongest $M_{2}$ tide [29, 40].
Figure 5: Pearson correlation coefficients between different models of the observed signals and a 3-D direct numerical simulation (denoted DNS) for the four tidal constituents. Correlations were calculated between real parts of the radial magnetic field maps at the altitude of 430 km (shown in Figures 2-3). Correlation coefficients for imaginary parts are similar.
@@
Although the structure of spectra are generally similar, there are some differences: SH degrees $\geq 15-20$ appear to be damped in the KALMAG model, which prevents comparison with other models at these degrees. The new GFO24 models exhibit better agreement with simulations at all scales compared to the predecessor study GO19. In particular, the low degree SH coefficients are closer to the 3-D simulations, highlighting a reduced contamination by external field sources. We stress that the presented SH power spectra were calculated at Earth’s surface. Such downward continuation is known to amplify noise [62], which reveals the intrinsic quality and resolution of the model without the effect of geometric attenuation of higher SH degrees. Strong noise levels, manifested by a growing spectrum with SH degree, or damping due to strong regularization/prior are more apparent when spectra are calculated at the surface. The generally good agreement, even for smaller $N_{2}$ and $O_{1}$ constituents, is a remarkable achievement of modern satellite magnetometry. That said, we do observe a larger misfit at higher SH degrees for the GFO24-HR models compared to the GFO24 models where they overlap, indicating higher variances at high SH degrees in the GFO24-HR models. This stems from a fundamental trade-off between the resolution and variance, which is inherent to the geomagnetic field inverse problem solved here.
Figure 6: The $R_{n}$ power spectra (eq. 5) of the time-averaged ocean tidal magnetic fields due to four tidal constituents for different models. Black lines denote 3-D numerical simulations (DNS). Spectra were calculated for the reference surface radius $R_{E}=6371.2$ km.